The function f: R→R defined by f(x)= | x-1|+ | 3-x| is differentiable at x= 4.



True or false with full explanation

1
Expert's answer
2022-03-01T09:42:16-0500
"f(x)= \begin{cases}
 -2x+4 &x<1 \\
 2 &1\le x<3 \\
2x-4 & x\ge 3
\end{cases}"

For "x\ge 3, f(x)=2x-4"

"f'(x)=2, x>3"

Therefore the function "f: R\u2192R" defined by "f(x)= | x-1|+ | 3-x|" is differentiable at "x= 4."